OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (anonymous):

For the real roots, you get factors of the form x-a. In this case, (x-5) and (x+3) are factors. In the case of the complex root, it works similarly, but we also need the conjugate. That means (x+1-3i) and (x+1+3i) are both factors. Multiply the four factors together, combine like terms, and you will have your answer.

OpenStudy (anonymous):

do we foil them together?

OpenStudy (anonymous):

\[(x-5)(x+3)(x+1-3i)(x+1+3i)=(x^2-2x-15)(x^2+2x+10)\]

OpenStudy (anonymous):

You should be able to finish from there.

OpenStudy (anonymous):

thank you! I understand it now

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

I did not find the right answer

OpenStudy (anonymous):

f(x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150 f(x) = x4 - 9x2 - 50x - 150

OpenStudy (anonymous):

those are my options

OpenStudy (anonymous):

@AnimalAin

OpenStudy (anonymous):

\[(x^2−2x−15)(x^2+2x+10)\]\[=x^2(x^2−2x−15)+2x(x^2−2x−15)+10(x^2−2x−15)\]\[=x^4−2x^3−15x^2+2x^3−4x^2−30x+10x^2−20x−150\]\[=x^4-9x^2-50x-150\]

OpenStudy (anonymous):

Looks like answer D. Make sure you pay attention to detail when you are working out your multiplication.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

do you know half angle rules for cos?

OpenStudy (anonymous):

In particular...?

OpenStudy (anonymous):

i figured it out

OpenStudy (anonymous):

Excellent. Do math every day.